The quality of the drinking water must be monitored as often as possible. One variable of concern is




the pH level, which measures the alkalinity or acidity of the water. A pH below 7.0 is acidic while a pH




above 7.0 is alkaline. A pH of 7.0 is neutral. A water-treatment plant is targeting higher than 8.0 pH. Based




on 16 random water samples, the mean and standard deviation were found to be: 𝑋 ̅=7.6 and s = 0.4. Test




the claim using 5%







Step




1 Describe the population parameter of interest




2 Formulate the null and alternative hypothesis




3 Check the assumptions




4 Choose a signifinance level size for α




5 Select the appropriate test statistic




Compute the test statistic using the appropriate formula




6 State the decision rule for rejecting or not the null hypothesis




7 Compare the computed test statistic and the critical value /s

1
Expert's answer
2022-05-27T15:02:52-0400

1. The pH level is the population parameter of interest


2.The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 8"

"H_1:\mu>8"


3.This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


4. Let the significance level be "\alpha = 0.05," "df=n-1=15" and the critical value for a right-tailed test is "t_c = 1.75305."

The rejection region for this right-tailed test is "R = \{t:t> 1.75305\}."


5. The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{7.6-8}{0.4/\sqrt{16}}=-4"

Since it is observed that "t=-4< 1.75305=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=15" degrees of freedom, "t=-4" is "p=0.99942," and since "p=0.99942>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 8, at the "\alpha = 0.05" significance level.